onsider rolling a fair 8 sided die. The sample space is S = {1, 2, 3, 4, 5, 6, 7, 8} Find the following probabilities: Round to 3 decimal places. a. P(2) =

b. P(odd number) =

c. P(not 7) =

d. P(less than 6) =

e. P(3.5) =
Answer #1

Sol)

Given

Sample space

S = { 1,2,3,4,5,6,7,8 }

Probability = favorable outcomes / total outcomes

Total outcomes = 8

a)

P( 2 ) =

There is only one 2 in the sample space

P( 2 ) = 1 / 8

=0.125

b)

P( odd number )

Odd numbers in sample space = ( 1,3,5,7 )

P( odd number ) = 4 / 8

= 0.5

c)

P( not 7 )

The numbers rather than 7 in sample space are

{ 1,2,3,4,5,6,8 }

P( not 7 ) = 7/8

= 0.875

d)

P( less than 6 )

The numbers less than 6 in sample space are

{ 1,2,3,4,5 }

P( less than 6 ) = 5/8

= 0.625

e)

P( 3.5 )

There is no number like 3.5 in the sample space

So probability

P( 3.5 ) = 0 / 8

= 0